# Knowledge Mechanisms in Large Language Models: A Survey and Perspective

### v1

@@ 1 Introduction
As we know, Large language models (LLMs) are also renowned for encapsulating extensive parametric knowledge Roberts et al. (2020); Sung et al. (2021); Cao et al. (2021); Zhong et al. (2021); Kandpal et al. (2023); Heinzerling and Inui (2020); Petroni et al. (2019); Qiao et al. (2023); Kritharoula et al. (2023); He et al. (2024a), achieving unprecedented progress in application. However, the knowledge mechanisms in LLMs for learning, storage, utilization, and evolution still remain mysterious Gould et al. (2023a). Extensive works aim to demystify various types of knowledge in LLMs through knowledge neurons Dai et al. (2022); Chen et al. (2024a) and circuits Elhage et al. (2021); Yao et al. (2024); Zou et al. (2024), yet these efforts, scattered across various tasks, await comprehensive review and analysis.
@@
As shown in Fig 1, this paper pioneeringly reviews the mechanism across the whole knowledge life cycle. We also propose a novel taxonomy for knowledge mechanisms in LLMs, as illustrated in Fig 2, which encompasses knowledge utilization at a specific time and knowledge evolution across all periods of LLMs  [^1] Knowledge utilization focuses on static knowledge at a specific period, while knowledge evolution explores the long-term dynamic development of knowledge across individual and group LLMs.. Specifically, we first introduce preliminaries of this field (§2) and review knowledge utilization mechanism from a new perspective (§3). Then, we delve into the fundamental principles for knowledge evolution (§4), discuss challenges of knowledge utilization, and posit some promising hypotheses to explore potential avenues for developing powerful and trustworthy models (§5). Finally, we also provide some future directions (§6) and tools for knowledge mechanism analysis (§C). Our contributions are as follows:
@@
fact knowledge Yu et al. (2023b); Jiang et al. (2024b), mental state Zhu et al. (2024) , utilization_leaf, text width=26.3em ] ] [ Connection, color=hidden-draw, fill=bluelight1!80, text width=4.3em, text=black [ 1) MLP + Attention: fact knowledge Geva et al. (2023); Lv et al. (2024); Li et al. (2023a)
@@
  • We suspect that the prevalent transformer architecture may impede creativity, data distribution and quantity may contribute to the fragility of parametric knowledge, leading hallucinations and knowledge conflict. Besides, dark knowledge will exist for a long time.
@@
Generally, this paper may help us to explore and manipulate advanced knowledge in LLMs, examine current limitations through the history of knowledge evolution, and inspire more efficient and trustworthy architecture and learning strategy for future models. Note that most hypotheses in this paper are derived from transformer-based LLMs. We also validate the generalizability of these hypotheses across other architectural models then propose universality intelligence in §B.
@@ 2.2 Definition of Knowledge in LLMs
$t\in\mathbf{T}\Rightarrow\mathcal{F}\ \text{masters knowledge}\ k,$
@@ 2.3 The Architecture of LLMs
The prevailing architecture in current LLMs is the Transformer. Specifically, a transformer-based LLM $\mathcal{F}$ begins with token embedding, followed by $L$ layers transformer block, and ends with token unembedding used for predicting answer tokens. Each block layer $l$ consists of Attention Heads (Attention) and Multilayer Perceptron (MLP):
@@
Other architectures including competitive variants of the transformer, e.g., SSM Gu and Dao (2023), TTT Sun et al. (2024) and RWKV Peng et al. (2023), and architectures in computer vision Li et al. (2023b) and multi-modal fields are detailed in §B.1.
@@ 2.4 Knowledge Analysis Methods
$t\leftarrow\mathcal{M}_{\mathbf{C}\subseteq\mathcal{F}}\left(r_{k\backslash t},\mathbf{C}\right),$ (3)
$t\in\mathbf{T}\Rightarrow\mathbf{C}\ \text{represents knowledge}\ k,$
@@
$t\leftarrow E_{\mathbf{C}\subseteq\mathcal{F}}\left(r_{k\backslash t},\mathbf{C},\mathcal{F}\right),$ (4)
$E$ is a evaluation metric, which can be a probe, logit lens, or a sparse representation. Probe is a meticulously trained classifier, and its classification performance is used to observe the relationship between model’s behavior and the output of $\mathbf{C}$ Räuker et al. (2023); Belinkov (2022); Elazar et al. (2021); McGrath et al. (2021); Gurnee et al. (2023). Logit lens usually translate output of $\mathbf{C}$ into vocabulary tokens via token unembedding nostalgebraist (2020); Geva et al. (2022b); Belrose et al. (2023); Pal et al. (2023); Din et al. (2024); Langedijk et al. (2023). Sparse representation disentangles the output of $\mathbf{C}$ into basic features via sparse autoencoder Gao et al. (2024b); Sharkey et al. (2022); Cunningham et al. (2023); Lee et al. (2006) or sparse dictionary learning He et al. (2024b); Olshausen and Field (1997); Yun et al. (2021). Each basic feature represents an interpretable unit, and the output of $\mathbf{C}$ is the combination Elhage et al. (2022); Bricken et al. (2023) of these basic features.
@@
$\mathbf{C}\leftarrow\mathcal{I}\left(r_{k\backslash t},\mathcal{F}\right),r\in\mathbf{R}_{k},$ (5)
$t\leftarrow E\left(r_{k\backslash t},\mathbf{C},\mathcal{F}\right)$
$\mathcal{I}$ is also known as causal mediation analysis Vig et al. (2020), causal tracing Meng et al. (2022), interchange interventions Geiger et al. (2022), activation patching Wang et al. (2023c); Zhang and Nanda (2023), and causal scrubbing techniques. For intervention-based methods, $E$ typically refers to token unembedding used for predicting answer tokens. Under evaluation metric of $E$, there exists a causal relationship between $\mathbf{C}$ and specific behavior of LLMs $\mathcal{F}$. Generally, $\mathcal{I}$ consists of the following three steps. 1) Clean run: perform an input with $\mathcal{F}$ and obtain the answer. 2) Corrupted run: corrupt the performing process by noisy input or activations of $\mathcal{F}$ Meng et al. (2022); Goldowsky-Dill et al. (2023); Stolfo et al. (2023); Yao et al. (2024); Conmy et al. (2023); Mossing et al. (2024); Lepori et al. (2023); Huang et al. (2023a). 3) Restoration run: recover the correct answer using $\mathbf{C}$ obtained from the above corrupted run Meng et al. (2022); Vig et al. (2020); Wang et al. (2023c); Zhang et al. (2017); Nanda (2023).
@@ 3 Knowledge Utilization in LLMs
According to Bloom’s Taxonomy of cognition levels Bloom et al. (1956); Keene et al. (2010); Fadul (2009), we categorize knowledge representation and utilization within LLMs into three levels (as shown in Fig 3): memorization, comprehension and application, and creation. Note that these mechanistic analyses are implemented via methods in §2.4. We further evaluate the applicability, advantages, and limitations of different methods in §3.4.
@@ 3.1 Memorization
Similar to the function regions of human brain Zhao et al. (2023a), this modular region hypothesis simplifies knowledge representation in transformer-based models into isolated modular region, e.g., MLPs or attention heads.
Knowledge is encoded via MLPs. Geva et al. (2021) posit that MLPs operate as key-value memories and each individual key vector corresponds to a specific semantic pattern or grammar. Based on the above finding, Geva et al. (2022b, a) reverse engineer the operation of the MLPs layers and find that MLPs can promote both semantic (e.g., measurement semantic including kg, percent, spread, total, yards, pounds, and hours) and syntactic (e.g., adverbs syntactic including largely, rapidly, effectively, previously, and normally) concepts in the vocabulary space. Miller and Neo (2024) find a single MLP neuron (in GPT-2 Large) capable of generating “an” or “a”. Subsequently, fact Dai et al. (2022); Meng et al. (2022) and commonsense knowledge Gupta et al. (2023) are found. Advanced language-specific neurons Tang et al. (2024), linguistic regions Zhao et al. (2023a), entropy neurons Stolfo et al. (2024), abstract conceptual Wang et al. (2024e) and unsafe Wang et al. (2024b); Wu et al. (2023) knowledge, are also observed in MLPs. In addition to MLP, knowledge is also conveyed by attention heads Geva et al. (2023); Gould et al. (2023b). Hoover et al. (2020) explain the knowledge each attention head has learned. Specifically, attention heads store evident linguistic features, positional information, and so on. Besides, fact knowledge Yu et al. (2023b); Li et al. (2023a) and bias Hoover et al. (2020) are mainly convey by attention heads. Jiang et al. (2024b) further observe that LLMs leverage self-attention to gather information through certain tokens in the contexts, which serve as clues, and use the value matrix for associative memory. Later, Zhu et al. (2024) also find that attention heads can simulate mental state and activate “Theory of Mind” (ToM) capability.
@@
Geva et al. (2023) outline the encoding of factual knowledge (e.g., “The capital of Ireland is Dublin”) through the following three steps: (1) subject (Ireland) information enrichment in MLPs, (2) the relation (capital of) propagates to the last token, (3) object (Dublin) is extracted by attention heads in later layers. This claim is supported by Li et al. (2024c). Similarly, Lv et al. (2024) conclude that task-specific attention head may move the topic entity to the final position of the residual stream, while MLPs conduct relation function.
Moreover, the recent prominent knowledge circuit framework Nainani (2024); Yao et al. (2024); He et al. (2024b); Elhage et al. (2021); Marks et al. (2024) advocates leveraging a critical computational subgraph among all components to explore internal knowledge within LLM parameters. The competencies for indirect object identification and color object tasks are discovered to be embedded in specialized knowledge circuits Conmy et al. (2023); Wang et al. (2023c); Merullo et al. (2023a); Yu et al. (2024c). Lan et al. (2024) also identify number-related circuits that encode the predictive ability of Arabic numerals, number words, and months. More importantly, experimental evidence demonstrates that various types of knowledge, including linguistic, commonsense, factual, and biased information, are encapsulated in specific knowledge circuits Yao et al. (2024). Interestingly, knowledge encoded by specific circuits can rival or even surpass that of the entire LLM. This may be because knowledge circuits memorized the relevant knowledge, while noise from other components might impede the model’s performance on these tasks.
@@ 3.3 Creation
However, from a statistical perspective, the intricate connections and activations between neurons, though not infinite, resist exhaustive enumeration. In terms of value, not all creations are valuable. Obtaining something valuable with an exceedingly low probability is impractical, as even a monkey could theoretically print Shakespeare’s works. How do LLMs ensure the probability of generating valuable creations? What are the mechanisms underlying the novelty and value of creation? A prevalent conjecture posits that novelty is generated through the random walk Sæbø and Brovold (2024). However, intuitively, current LLMs themselves seem unable to evaluate the value of creations due to architectural limitations Chakrabarty et al. (2024). Because, once the next token is generated, there is no intrinsic mechanism for accepting or rejecting the creations. This hinders the evaluation of the usefulness and value of proposed novelties, as humans do, by bending, blending, or breaking biases Sæbø and Brovold (2024). Some works assume that each token is indeed valuable and meets long-term expectations. However, the well-known hallucination problem Xu et al. (2024d) of LLMs refutes this assumption. Besides, the transformer architecture struggles with long context Li et al. (2024a), despite the existence of many variants for addressing this issue. More importantly, MLPs of Transformer may also work contrary to creativity, i.e., the increased attentions narrow the conditional distribution for token prediction Sæbø and Brovold (2024).
@@ 4.1 Individual Evolution
After pre-training, LLMs are anticipated to refresh their internal knowledge to keep pace with the evolving world during post-training stage. Although LLMs seem to absorb new knowledge through continued learning, follow user instructions via instruct tuning Zhang et al. (2023c), and align with human values through alignment tuning Ziegler et al. (2019), Ji et al. (2024a) have noted that LLMs intrinsically resist alignment during the post-training phase. In other words, LLMs tend to learn factual knowledge through pre-training, whereas fine-tuning  [^3] Fine-tuning includes instruct tuning and alignment tuning Zhao et al. (2023b). teaches them to utilize it more efficiently Gekhman et al. (2024); Zhou et al. (2023a); Ovadia et al. (2024). Ren et al. (2024) also posit that instruction tuning is a form of self-alignment with existing internal knowledge rather than a process of learning new information. We conjecture that the debate on whether these processes truly introduce new knowledge stems from information conflicts. For example, the conflict between outdated information within LLMs and new external knowledge exacerbates their difficulty in learning new information. To mitigate information conflicts, Ni et al. (2023) propose first forgetting old knowledge then learning new knowledge. Another technique, retrieval-augmented generation (RAG) Huang and Huang (2024), while avoiding conflicts within internal parameters, still needs to manage conflicts between retrieved external information and LLMs’ internal knowledge Xu et al. (2024b). RAG also attempt to efficiently and effectively integrate new knowledge across passages or documents using multiple retrieval Yang et al. (2024a) and hippocampal indexing Gutiérrez et al. (2024). Besides, editing technologies, including knowledge and representation editing, exhibit promising potential for knowledge addition, modification, and erasure. Specifically, knowledge editing Zhang et al. (2024a); Wang et al. (2023d); Mazzia et al. (2023) aims to selectively modify model parameters responsible for specific knowledge retention, while representation editing Zou et al. (2023); Wu et al. (2024) adjusts the model’s conceptualization of knowledge to revise the stored knowledge within LLMs. Note that the other strategy for knowledge editing adds external parameters or memory banks for new knowledge while preserving models’ parameters. We also provide the comparison of the above methods in §A for better understanding.

### v2

@@ 1 Introduction
As we know, Large language models (LLMs) are renowned for encapsulating extensive parametric knowledge Roberts et al. (2020); Sung et al. (2021); Cao et al. (2021a); Zhong et al. (2021); Kandpal et al. (2023); Heinzerling and Inui (2020); Petroni et al. (2019); Qiao et al. (2023); Kritharoula et al. (2023); He et al. (2024a), achieving unprecedented progress in application. However, the knowledge mechanisms in LLMs for learning, storage, utilization, and evolution still remain mysterious Gould et al. (2023a). Extensive works aim to demystify various types of knowledge in LLMs through knowledge neurons Dai et al. (2022); Chen et al. (2024a) and circuits Elhage et al. (2021); Yao et al. (2024); Zou et al. (2024), yet these efforts, scattered across various tasks, await comprehensive review and analysis.
@@
As shown in Fig 1, this paper pioneeringly reviews the mechanism across the whole knowledge life cycle. We also propose a novel taxonomy for knowledge mechanisms in LLMs, as illustrated in Fig 2, which encompasses knowledge utilization at a specific time and knowledge evolution across all periods of LLMs  [^1] Knowledge utilization focuses on static knowledge at a specific period, while knowledge evolution explores the long-term dynamic development of knowledge across individual and group LLMs.. Specifically, we introduce preliminaries of this field (§2) and review the knowledge utilization mechanism from a new perspective (§3), delve into the fundamental principles for knowledge evolution (§4). Then, we investigate how to construct more efficient and trustworthy LLMs from the perspective of knowledge mechanism (§5). Later, We discuss open questions about the knowledge LLMs have and have not acquired (§6). Finally, we also provide some future directions (§7) and tools for knowledge mechanism analysis (§C). Our contributions are as follows:
@@
fact knowledge Yu et al. (2023c); Jiang et al. (2024b), mental state Zhu et al. (2024) , utilization_leaf, text width=26.3em ] ] [ Connection, color=hidden-draw, fill=bluelight1!80, text width=4.3em, text=black [ 1) MLP + Attention: fact knowledge Geva et al. (2023); Lv et al. (2024); Li et al. (2023a)
@@
  • We observe that LLMs have learned basic world knowledge. However, the learned knowledge is fragile, leading to challenges such as hallucinations and knowledge conflicts. We speculate that this fragility may be primarily due to improper learning data. Besides, the unlearned dark knowledge will exist long.
@@
Generally, this paper may help us to explore and manipulate advanced knowledge in LLMs, examine current limitations through the history of knowledge evolution, and inspire more efficient and trustworthy architecture and learning strategy for future models from knowledge mechanism perspective. Note that most hypotheses in this paper are derived from transformer-based LLMs. We also validate the generalizability of these hypotheses across other architectural models and then propose universality intelligence in §B.
@@ 2.2 Definition of Knowledge in LLMs
$(t\in\mathbf{T})\Rightarrow(\mathcal{F}\ \text{masters knowledge}\ k),$
@@ 2.3 The Architecture of LLMs
The prevailing architecture in current LLMs is the Transformer Vaswani et al. (2017). Specifically, a transformer-based LLM $\mathcal{F}$ begins with token embedding, followed by $L$ layers transformer block, and ends with token unembedding used for predicting answer tokens. Each block layer $l$ consists of Attention Heads (Attention) and Multilayer Perceptron (MLP):
@@
Other architectures including competitive variants of the transformer, e.g., SSM Gu and Dao (2023), TTT Sun et al. (2024) and RWKV Peng et al. (2023), and architectures in computer vision Li et al. (2023c) and multi-modal fields are detailed in §B.1.
@@ 2.4 Knowledge Analysis Methods
$t=\mathcal{M}_{\mathbf{C}\subseteq\mathcal{F}}\left(r_{k\backslash t},\mathbf{C}\right),$ (3)
$(t\in\mathbf{T})\Rightarrow(\mathbf{C}\ \text{represents knowledge}\ k),$
@@
$t=E_{\mathbf{C}\subseteq\mathcal{F}}\left(r_{k\backslash t},\mathbf{C},\mathcal{F}\right),$ (4)
$E$ is a evaluation metric, which can be a probe Räuker et al. (2023), logit lens nostalgebraist (2020), or a sparse representation Gao et al. (2024b). Probe is a meticulously trained classifier, and its classification performance is used to observe the relationship between model’s behavior and the output of $\mathbf{C}$ Belinkov (2022); Elazar et al. (2021); McGrath et al. (2021); Gurnee et al. (2023). Logit lens usually translate output of $\mathbf{C}$ into vocabulary tokens via token unembedding Geva et al. (2022b); Belrose et al. (2023); Pal et al. (2023); Din et al. (2024); Langedijk et al. (2023). Sparse representation disentangles the output of $\mathbf{C}$ into basic features via sparse autoencoder Sharkey et al. (2022); Cunningham et al. (2023); Lee et al. (2006) or sparse dictionary learning He et al. (2024b); Olshausen and Field (1997); Yun et al. (2021). Each basic feature represents an interpretable unit, and the output of $\mathbf{C}$ is the combination Elhage et al. (2022); Bricken et al. (2023) of these basic features.
@@
$\mathbf{C}=\mathcal{I}\left(r_{k\backslash t},\mathcal{F}\right),$ (5)
$t=E\left(r_{k\backslash t},\mathbf{C},\mathcal{F}\right)$
$\mathcal{I}$ is also known as causal mediation analysis Vig et al. (2020), causal tracing Meng et al. (2022), interchange interventions Geiger et al. (2022), activation patching Wang et al. (2023c); Zhang and Nanda (2023), path patching Goldowsky-Dill et al. (2023), and causal scrubbing techniques LawrenceC et al. (2022). Specifically, $\mathcal{I}$ consists of the following three steps. 1) Clean run: $\mathcal{F}$ generates the correct answer $t$ based on the input $r_{k\backslash t}$. 2) Corrupted run: corrupt the generation process of $\mathcal{F}$ in the clean run by introducing noise into the input or neurons Meng et al. (2022); Goldowsky-Dill et al. (2023); Stolfo et al. (2023); Yao et al. (2024); Conmy et al. (2023); Mossing et al. (2024); Lepori et al. (2023); Huang et al. (2023a). 3) Restoration run: recover the correct answer $t$ by restoring unnoised information from $\mathbf{C}$ Meng et al. (2022); Vig et al. (2020); Wang et al. (2023c); Zhang et al. (2017); Nanda (2023). For intervention-based methods, $E$ typically refers to the token unembedding used for predicting answer tokens. Under the evaluation metric $E$, there exists a causal relationship between $\mathbf{C}$ and specific behavior of LLMs $\mathcal{F}$ in Eq 5.
@@ 3 Knowledge Utilization in LLMs
Inspried by Bloom’s Taxonomy of cognition levels Wilson (2016); Bloom et al. (1956); Keene et al. (2010); Fadul (2009), we categorize knowledge representation and utilization within LLMs into three levels (as shown in Fig 3): memorization, comprehension and application, and creation 333 Note that we combine analyzing, evaluating, and creating from Bloom’s Taxonomy into one category level (creation) in our taxonomy, as they are difficult to disentangle. Specifically, creation emphasizes the capacity and process of forming novel and valuable things. Analyzing Wilson (2016), which breaks materials or concepts into parts, is used for creating novel things. Evaluating Wilson (2016) is usually used for assessing the value of new creations.. Note that these mechanistic analyses are implemented via methods in §2.4. We further evaluate the applicability, advantages, and limitations of different methods in §3.4.
@@ 3.1 Memorization
This modular region hypothesis simplifies knowledge representation in transformer-based models into isolated modular region, e.g., MLPs or attention heads. Knowledge is encoded via MLPs. Geva et al. (2021) posit that MLPs operate as key-value memories and each individual key vector corresponds to a specific semantic pattern or grammar. Based on the above finding, Geva et al. (2022b, a) reverse engineer the operation of the MLPs layers and find that MLPs can promote both semantic (e.g., measurement semantic including kg, percent, spread, total, yards, pounds, and hours) and syntactic (e.g., adverbs syntactic including largely, rapidly, effectively, previously, and normally) concepts in the vocabulary space. Miller and Neo (2024) find a single MLP neuron (in GPT-2 Large) capable of generating “an” or “a”. Subsequently, fact Dai et al. (2022); Meng et al. (2022) and commonsense knowledge Gupta et al. (2023) are found. Advanced language-specific neurons Tang et al. (2024), linguistic regions Zhao et al. (2023a), entropy neurons Stolfo et al. (2024), abstract conceptual Wang et al. (2024e) and unsafe Wang et al. (2024b); Wu et al. (2023) knowledge, are also observed in MLPs. In addition to MLP, knowledge is also conveyed by attention heads Geva et al. (2023); Gould et al. (2023b). Hoover et al. (2020) explain the knowledge each attention head has learned. Specifically, attention heads store evident linguistic features, positional information, and so on. Besides, fact knowledge Yu et al. (2023c); Li et al. (2023a) and bias Hoover et al. (2020) are mainly convey by attention heads. Jiang et al. (2024b) further observe that LLMs leverage self-attention to gather information through certain tokens in the contexts, which serve as clues, and use the value matrix for associative memory. Later, Zhu et al. (2024) also find that attention heads can simulate mental state and activate “Theory of Mind” (ToM) capability.
@@
Geva et al. (2023) outline the encoding of factual knowledge (e.g., “The capital of Ireland is Dublin”) through the following three steps: (1) subject (Ireland) information enrichment in MLPs, (2) the relation (capital of) propagates to the last token, (3) object (Dublin) is extracted by attention heads in later layers. This claim is supported by Li et al. (2024c). Similarly, Lv et al. (2024) conclude that task-specific attention head may move the topic entity to the final position of the residual stream, while MLPs conduct relation function. Moreover, the recent prominent knowledge circuit framework Nainani (2024); Yao et al. (2024); He et al. (2024b); Elhage et al. (2021); Marks et al. (2024) advocates leveraging a critical computational subgraph among all components to explore internal knowledge within LLM parameters. The competencies for indirect object identification and color object tasks are discovered to be embedded in specialized knowledge circuits Conmy et al. (2023); Wang et al. (2023c); Merullo et al. (2023a); Yu et al. (2024c). Lan et al. (2024) also identify number-related circuits that encode the predictive ability of Arabic numerals, number words, and months. More importantly, experimental evidence demonstrates that various types of knowledge, including linguistic, commonsense, factual, and biased information, are encapsulated in specific knowledge circuits Yao et al. (2024). Interestingly, knowledge encoded by specific circuits can rival or even surpass that of the entire LLM. This may be because knowledge circuits memorized the relevant knowledge, while noise from other components might impede the model’s performance on these tasks.
@@ 3.3 Creation
However, from a statistical perspective, the intricate connections and activations between neurons, though not infinite, resist exhaustive enumeration. In terms of value, not all creations are valuable. Obtaining something valuable with an exceedingly low probability is impractical, as even a monkey could theoretically print Shakespeare’s works. How do LLMs ensure the probability of generating valuable creations? What are the mechanisms underlying the novelty and value of creation? A prevalent conjecture posits that novelty is generated through the random walk Sæbø and Brovold (2024). However, intuitively, current LLMs themselves seem unable to evaluate the value of creations due to architectural limitations Chakrabarty et al. (2024). Because, once the next token is generated, there is no intrinsic mechanism for accepting or rejecting the creations. This hinders the evaluation of the usefulness and value of proposed novelties, as humans do, by bending, blending, or breaking biases Sæbø and Brovold (2024). Some works assume that each token is indeed valuable and meets long-term expectations. However, the well-known hallucination problem Xu et al. (2024d) of LLMs refutes this assumption. Besides, the transformer architecture struggles with long context Li et al. (2024a), despite the existence of many variants for addressing this issue Huang et al. (2023c); Liu et al. (2024b); Zhang et al. (2024b); Fountas et al. (2024). More importantly, MLPs of Transformer may also work contrary to creativity, i.e., the increased attentions narrow the conditional distribution for token prediction Sæbø and Brovold (2024).
@@ 4.1 Individual Evolution
After pre-training, LLMs are anticipated to refresh their internal knowledge to keep pace with the evolving world during post-training stage. Although LLMs seem to absorb new knowledge through continued learning, follow user instructions via instruct tuning Zhang et al. (2023c), and align with human values through alignment tuning Ziegler et al. (2019), Ji et al. (2024a) have noted that LLMs intrinsically resist alignment during the post-training phase. In other words, LLMs tend to learn factual knowledge through pre-training, whereas fine-tuning  [^4] Fine-tuning includes instruct tuning and alignment tuning Zhao et al. (2023b). teaches them to utilize it more efficiently Gekhman et al. (2024); Zhou et al. (2023a); Ovadia et al. (2024). Ren et al. (2024) also posit that instruction tuning is a form of self-alignment with existing internal knowledge rather than a process of learning new information. We conjecture that the debate on whether these processes truly introduce new knowledge stems from information conflicts. For example, the conflict between outdated information within LLMs and new external knowledge exacerbates their difficulty in learning new information. To mitigate information conflicts, Ni et al. (2023) propose first forgetting old knowledge then learning new knowledge. Another technique, retrieval-augmented generation (RAG) Huang and Huang (2024), while avoiding conflicts within internal parameters, still needs to manage conflicts between retrieved external information and LLMs’ internal knowledge Xu et al. (2024b). RAG also attempt to efficiently and effectively integrate new knowledge across passages or documents using multiple retrieval Yang et al. (2024a) and hippocampal indexing Gutiérrez et al. (2024). Besides, editing technologies, including knowledge and representation editing, exhibit promising potential for knowledge addition, modification, and erasure. Specifically, knowledge editing Meng et al. (2022); Mitchell et al. (2022); Cao et al. (2021b); Zhang et al. (2024a); Wang et al. (2023d); Mazzia et al. (2023) aims to selectively modify model parameters responsible for specific knowledge retention, while representation editing Zou et al. (2023); Wu et al. (2024) adjusts the model’s conceptualization of knowledge to revise the stored knowledge within LLMs. Note that the other strategy for knowledge editing adds external parameters or memory banks for new knowledge while preserving models’ parameters. We also provide the comparison of the above methods in §A for better understanding.